OpenStudy (anonymous):

A company policy requires that, for every 30 employees, there must be 3 supervisors. If there are 273 supervisors at the company, how many employees does the company have a.$60.80 b.$129.64 c.$139.96 d.$69.60

OpenStudy (texaschic101):

30/3 = x/273 ---( 30 employees to 3 supervisors = x employees to 273 supervisors) this is a proportion, so we now cross multiply 3x = (273)(30) 3x = 8190 x = 8190/3 x = 2730 there are 2730 employees when there is 273 supervisors

OpenStudy (texaschic101):

whats with the money answer choices ??

OpenStudy (anonymous):

sorry lol I put the other thing in so sorry